A silver block has a mass of 0.337 kg and is initially at a temperature of 18.4° C. How much energy must be transferred to the block as heat for it to reach a final temperature of 44.2° ;C? Use csilver =235 J/kg•K for the specific heat value of silver.

  • -3,500 J
  • 3,500 J
  • -1,460 J
  • 2,040 J
  • -2,040 J
  • 1,460 J
Answer #1

T, = 18.4°C, Tz = 44.2°C, C = 235 g/kg.K m=0.337 kg Q = MCAT = (0.337) (235) (44.2-18.4) = 2043.2) => T0 = 2040 joule is clos